OpenStudy (anonymous):

y'=sqrt(x+y+1) using y=vt substitution but I can't seem to find a way to separate the variables

OpenStudy (anonymous):

\[y'=\sqrt{x+y+1}\] becomes: \[v+x \frac{ dv }{ dx } = \sqrt{x+vx+1}\]

OpenStudy (anonymous):

Ehh ill answer this one haha just give me a second!

OpenStudy (anonymous):

thanks been chasing my tail for an hour

OpenStudy (anonymous):

Lets see hrmmmm...lol just give me a sec to finish some other calculus loose ends elsewhere

OpenStudy (anonymous):

Hello! Back again haha

OpenStudy (anonymous):

Lets see...

OpenStudy (anonymous):

Well, I would divide by \(x\) to get rid of the \(x\) attached to the differential first!

OpenStudy (anonymous):

What have you done so far?

OpenStudy (anonymous):

oh sorry I wandered off too...trying to find a bernoulli tutorial anyway what haven't I tried

OpenStudy (anonymous):